A carnot engine takes 3 × 106 cal of heat from a reservoir at 627°C, and gives it to a sink at 27°C. The work done by the engine is

(A) 4.2 × 106 J

(B) 8.4 × 106 J

(C) 16.8 × 106 J

(D) zero

Grade:upto college level

1 Answers

jitendra
6 years ago
(b) 8.4*106
work done by enigne while taking heat 
Q=3*106 cal is w=2*106*4.2=8.4*106 j.
this is expect answer i think